OpenStudy (anonymous):

Solve the equation by making an appropriate substitution. x^4 + 2x^2 - 24 = 0

OpenStudy (ybarrap):

let u = x^2 then solve u^2 + 2u - 24 =0 (u+6)(u-4)=0 substitute x^2 for u to complete Make sense?